OpenStudy (anonymous):

Write An equation Perpendicular to the given equation through the ordered pair y=4x+6(1,-3). Please Show Steps,Instruction on how to solve this ..Thanxx

OpenStudy (campbell_st):

perpendicular to the given like has a gradient of -1/4 as the product of the gradients is -1 then using point gradient formula y + 3 = -1/4(x - 1) 4y + 3 = -x + 1 or x + 4y + 2 = 0 or y = -1/4 x -2

OpenStudy (anonymous):

The general formula of a straight line is: y = mx + c Where m is the gradient and c is the y-intercept. So the gradient of the line you have been given is 4. This means the gradient of a line perpendicular to it will have a gradient of "-1/4" You can find this by = this equation gradient of eq1 * gradient of line perpendicular = -1 so \[4 \times x = -1\] \[x= -1/4 \] So the line will have the equation: y = -1/4x + c where c is the y-intercept then u have to use the gradient formular (y-y1) = m(x-x1) \[y-(-3)=\frac{1}{4}(x-(1))\] \[y+3=-\frac{1}{4}x +\frac{1}{4}\] y= -1/4x-11/4